determine the point of discontinuity of the function f and the nature of discontinuity at each points

f(x)= { -x^2, when x<=0

{ 4-5x, when 0<x<=1

{ 3x-4x^2 when 1<x<=2

{-12x+2x when x>2

and check whether function f is derivable at x=1


1
Expert's answer
2021-05-07T09:02:22-0400
"f(x) = \begin{cases}
 -x^2 &\text{when } x\leq0 \\
 4-5x &\text{when } 0<x\leq1\\
 3x-4x^2 &\text{when } 1<x\leq2\\
 -12x+2x &\text{when } x>2\\
\end{cases}"


"\lim\limits_{x\to0^-}f(x)=\lim\limits_{x\to0^-}(-x^2)=0"

"\lim\limits_{x\to0^+}f(x)=\lim\limits_{x\to0^+}(4-5x)=4"

"\lim\limits_{x\to0^-}f(x)=0\not=4=\lim\limits_{x\to0^+}f(x)"

"\lim\limits_{x\to0}f(x)" does not exist.

The function "f(x)" is discontinuous at "x=0" and has a jump discontinuity at "x=0."



"\lim\limits_{x\to1^-}f(x)=\lim\limits_{x\to1^-}(4-5x)=-1"

"\lim\limits_{x\to1^+}f(x)=\lim\limits_{x\to1^+}(3x-4x^2)=-1"

"\lim\limits_{x\to1^-}f(x)=-1=\lim\limits_{x\to1^+}f(x)"

"\lim\limits_{x\to1}f(x)=-1".


"f(1)=4-5(1)=-1"

The function "f(x)" is continuous at "x=1."



"\lim\limits_{x\to2^-}f(x)=\lim\limits_{x\to2^-}(3x-4x^2)=-10"

"\lim\limits_{x\to2^+}f(x)=\lim\limits_{x\to2^+}(-12x+2x)=-20"

"\lim\limits_{x\to2^-}f(x)=-10\not=-20=\lim\limits_{x\to2^+}f(x)"

"\lim\limits_{x\to2}f(x)" does not exist.

The function "f(x)" is discontinuous at "x=2" and has a jump discontinuity at "x=2."



"f(1)=-1"


"\lim\limits_{h\to0^-}\dfrac{f(1+h)-f(1)}{h}=\lim\limits_{h\to0^-}\dfrac{4-5(1+h)-(-1)}{h}"

"=\lim\limits_{h\to0^-}\dfrac{-5h}{h}=-5"

"\lim\limits_{h\to0^+}\dfrac{f(1+h)-f(1)}{h}"

"=\lim\limits_{h\to0^+}\dfrac{3(1+h)-4(1+h)^2-(-1)}{h}"

"=\lim\limits_{h\to0^+}\dfrac{3+3h-4-8h-4h^2+1}{h}"

"=\lim\limits_{h\to0^+}(-5-4h)=-5"


"\lim\limits_{h\to0^-}\dfrac{f(1+h)-f(1)}{h}=-5"

"=\lim\limits_{h\to0^+}\dfrac{f(1+h)-f(1)}{h}"

Then


"\lim\limits_{h\to0}\dfrac{f(1+h)-f(1)}{h}=-5"

The function "f(x)" is derivable at "x=1," and


"f'(1)=-5"